Determine the amplitude, period, and phase shift of the function. Graph the function. y= sin (4x -21)
Answer #1

y = sin(4x-2x) - Sin(4(2-1/2)) to Y = A Sin(B1%-C))+d Amplitude = 1 period = 20 = 27 = 1/2 phaseshift = 1/2 Vetical Shift - 0